OpenStudy (anonymous):

How many x-intercepts does the graph of y = x2 have?

OpenStudy (anonymous):

0 1 2 3 are the choices

hero (hero):

Hint: Graph y = x^2

OpenStudy (anonymous):

how?

hero (hero):

Graphing Calc

OpenStudy (cwrw238):

when y=0 what is x?

OpenStudy (anonymous):

im doing EOC practice (EOC is like a standardized test) and we dont use graphing calcs on it

OpenStudy (anonymous):

@cwrw238 there is no x in that statement...

OpenStudy (anonymous):

x-intercept is when y=0; So solve for 0=x2

OpenStudy (cwrw238):

y = x^2!!!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

0?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

so 0,0 is a point where the graph passes through are there any others?

OpenStudy (anonymous):

i dont think so.... so its B?

OpenStudy (cwrw238):

x^2 = 0 just has one root right?

OpenStudy (anonymous):

yeah

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

so its B because it ONLY goes through 0,0?

OpenStudy (cwrw238):

the graph just touches 0,0 

OpenStudy (anonymous):

oh ok thanks!

OpenStudy (cwrw238):

yw